# Using VLM Reasoning to Constrain Task and Motion Planning

### v1

@@ VII Conclusion
We have presented VIZ-COAST, a TAMP solver that uses VLMs’ common-sense geometric reasoning capabilities to constrain TAMP problems a priori. Unlike prior works, the constraints produced by the VLM generalize to unseen problem instances within a broader domain and incur minimal additional overhead by directly interfacing through APIs to an SMT-based task planner. These constraints eliminate potential downward refinement failures that arise between the symbolic abstraction and geometry of the world—preventing the planner from attempting to pick obstructed objects, place objects in cluttered locations, and perform other infeasible actions. We have demonstrated that VIZ-COAST effectively leverages VLMs’ insights to improve planning times by reducing or even eliminating replanning due to downward refinement failures. For example, in the Blocks domain, VIZ-COAST produced general constraints on the first attempt, resulting in task plans that were refinable on the first attempt for all 35 problem instances. In the Containers domain, VIZ-COAST produces constraints that enable solutions to all 30 instances, while baseline methods fail to find solutions due to the complexity of the unconstrained search space.
Overall, VIZ-COAST demonstrates that VLMs can provide valuable insights into problem structure that accelerate search performance. These results are promising for future applications that extract even richer structural information, operate in online settings, or require constraint insertion based on natural language instructions grounded to the physical world. One limitation of our approach is that VLM inference introduces unpredictable latencies. Our method is agnostic to the specific VLM model used; as models become faster and more efficient, we expect this variance to decrease. Evaluating our methodology with local or distilled models represents an interesting direction for future work. We also evaluated our approach exclusively in simulation environments. While we believe that the generality of large pre-trained VLMs will enable constraint extraction in real-world settings, empirical validation of this capability remains future work.

### v2

@@ VII Conclusion
We have presented VIZ-COAST, a TAMP approach that uses VLMs’ common-sense geometric reasoning capabilities to constrain TAMP problems. Unlike prior works, the constraints produced generalize to unseen problem instances within a broader domain and incur minimal overhead by directly interfacing with an SMT-based task planner. These constraints prevent downward refinement failures when grounding the symbolic task plan to continuous motion. Furthermore, VIZ-COAST’s closed-loop mechanism for updating constraints with planner feedback enables reasoning over several problem instances to solve difficult domains. We have demonstrated that our approach improves planning times and success rates by reducing downward refinement failures and eschewing the need for auxiliary symbolic constructions required by baselines. For example, in the Blocks and Containers domains, VIZ-COAST produced general constraints on the first attempt, resulting in task plans that were refinable within $K$ incremental attempts for all 30 problem instances. In the Containers and Packing domains, VIZ-COAST solves all problems, while baseline methods often fail to find solutions due to the complexity introduced by auxiliary constructions. Finally, we empirically demonstrate that the performance of VIZ-COAST is robust to the VLM model used, and that current standard services are fast and powerful enough to be practical through our method.
Overall, VIZ-COAST demonstrates that VLMs can provide valuable insights into problem structure that accelerate task-and-motion-planning. These results are promising for future applications that may plan using richer abstraction, operate in dynamic online settings, or require constraint insertion based on natural language instructions grounded to the physical world. Although we have validated the effectiveness of VLM models across several providers, our reliance on a proprietary cloud-based service remains a source of variance. Evaluating our method with local or distilled models is an interesting direction for future work. Additionally, the generalization capacity of inferred constraints is currently limited to problems within the same PDDL domain. Future work could explore transferring or composing constraints learned from simpler domains to solve more complex ones. Lastly, we evaluated our approach exclusively in simulation. While we believe that the generality of large pre-trained VLMs will enable constraint synthesis in a wide range of settings, empirical validation of VIZ-COAST’s ability to handle the geometric and perceptual complexity of the real world remains future work.